OpenStudy (31356):

Distributive Property V= 36(a-c)-24c(2a-3c)+2a(-2(3a+c) a= 7 c=9 MEDAL REWARDED!!!

OpenStudy (31356):

@texaschic101 @mathslover @whpalmer

OpenStudy (31356):

@whpalmer4

OpenStudy (31356):

Please help o.0

OpenStudy (texaschic101):

V = 36(a - c) - 24c(2a - 3c) + 2a(-2(3a + c)) a = 7 c = 9 V = 36(7 - 9) - 24(9)(2(7) - 3(9)) + 2(7)(-2(3(7) + 9) V = 36(-2) - 216(14 - 27) + 14(-2(30)) V = - 72 - 216(-13) + 14(-60) V = -72 + 2898 - 840 V = - 912 + 2898 V = 1986

OpenStudy (texaschic101):

oops....hold on V = -72 - 216(-13) + 14(-60) V = -72 + 2808 - 840 V = 2808 - 912 V = 1896 sorry about that :(

OpenStudy (31356):

That's okay thanks though :D